solve for x for the gven equation of inverse trigonometric function:
arcsin(x) + arcsin(1-x) = arccos(x)?
 

Grade:12

1 Answers

Lab Bhattacharjee
8 years ago
\text{ HINT: Set} \arcsin x= y\implies x=\sin y\\\text{and use}\arcsin x+\arccos x=\dfrac\pi2 \\ \text{ then apply } \sin \text{ on both side to find } x=0,\dfrac12